OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction.

OpenStudy (jdoe0001):

what's the slope of the y = 4x +2?

OpenStudy (kropot72):

The slope of the line y = 4x + 2 is 4. Let the unknown slope = m The product of the two slopes must equal -1 if the two lines are perpendicular. 4m = -1 .............(1) Can you solve equation (1) to find the value of m ?

OpenStudy (anonymous):

im not sure

OpenStudy (jdoe0001):

y = 4x + 2 ^ slope as kropot72 already pointed out

OpenStudy (anonymous):

Formula is y=mx+c so here the question is y = 4x + 2 so 4 is the slope :)

OpenStudy (kropot72):

4m = -1 What do you get if you divide both sides by 4 ?

OpenStudy (anonymous):

-4

OpenStudy (kropot72):

What is 4m/4 ? \[\frac{4m}{4}=?\]

OpenStudy (anonymous):

1

OpenStudy (kropot72):

No. \[\frac{4}{4}=1\] \[\frac{4m}{4}=?\]

OpenStudy (anonymous):

m?

OpenStudy (jdoe0001):

perpendicular lines, have "negative reciprocal slopes", so if one function has a slope of "4" reciprocal of 4 = 1/4 negativize that = -(1/4) so, that'd be the negative reciprocal of "4", and thus the slope of the 1st function which passes through (4, -2)

OpenStudy (anonymous):

so its 1/4?

OpenStudy (anonymous):

or -1/4 i mean

OpenStudy (jdoe0001):

or 4m = -1 => m = -1/4 as kropot72 indicated

OpenStudy (anonymous):

ok thanks !

OpenStudy (jdoe0001):

now, use your slope and the points it passes through to get the function, using the point-slope form http://www.teachhandson.com/calculationb.gif